OpenStudy (anonymous):

Which of the relations has a domain of {-5, 0, 5}? {(-5, 0), (0, 5), (1, 0)} {(-5, 5), (0, 5), (5,5)} {(0, -5), (0, 0), (0, 5)}

pooja195 (pooja195):

Domain=x set of values

pooja195 (pooja195):

Match the x values (given) with the options

jimthompson5910 (jim_thompson5910):

Example: Let's say we're given this relation (set of ordered pairs) \[\Large \left\{(1,2),(5,7),(88,12)\right\}\] Let's mark the x coordinate of each point in red \[\Large \left\{({\color{red}{1}},2),({\color{red}{5}},7),({\color{red}{88}},12)\right\}\] as @pooja195 is saying, the domain is the set of x values of each point. So in my example, the domain is this set \[\Large \left\{{\color{red}{1}},{\color{red}{5}},{\color{red}{88}}\right\}\]